OpenStudy (anonymous):

a) write a inequality that presents the interval b) State whether the interval bouded or unbouded . [0,9)

OpenStudy (zarkon):

you want this... \[0\leq x<9\]

OpenStudy (anonymous):

a or b?

OpenStudy (zarkon):

a

OpenStudy (anonymous):

b?

OpenStudy (zarkon):

it is bounded

OpenStudy (anonymous):

what is mean bounded or unbounded?

OpenStudy (zarkon):

bounded...means bounded above and below.. for bounded above..there exists a number y such that for all x in your interval \[x\leq y\]

OpenStudy (zarkon):

for bounded below..there exists a number y such that for all x in your interval \[x\geq y\]

OpenStudy (zarkon):

an example of an unbounded interval is \[[5,\infty)\]

OpenStudy (anonymous):

how about [-1,5] the same ?

OpenStudy (zarkon):

\[-1\leq x\leq 5\] it is bounded

OpenStudy (anonymous):

\[(11,\infty) = unbouded\]\[(-\infty,-2)= unbouded\]?

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

ty